OpenStudy (anonymous):

When attempting to do his/her homework, a student uses the addition property of equality to solve for a below. Which statement BEST applies to the sample below? a + 3 = 9 a + 3 – 3 = 9 – 9 a = 0 A. The student failed to keep both sides of the equation balanced. B. The student’s final answer does not have the variable a isolated. C. The use of the addition property of equality is sound, and the answer is correct. D. The student should have added 3 to the left side of the equation rather than subtracting.

OpenStudy (anonymous):

Ehm. is there an error in the way to solve the equation?

OpenStudy (anonymous):

No that is how the question is written

OpenStudy (anonymous):

No, i mean here is how they show you the supposed method: a+3=9 a+3+(-3)=9+(-9) Is that okay? (With your mathematical knowledge)

OpenStudy (anonymous):

No I find it invalid for them to do that. I was confused like crazy that is why I asked for help

OpenStudy (anonymous):

Yeah, exactly, is invalid. Now, the question is which answer describes the situation at best. Analyze each case. Then you'll see that only one matches the problem.

OpenStudy (anonymous):

Ok thank you Umangiased

OpenStudy (anonymous):

No problem ;)